# The Illusive Slump of Disruptive Patents

### v1

@@ B CD Index and Patent Law Change
Figure B.1: Excluding Backward Citations to Patent Applications Affects the CD Index
@@
First, as Fig. B.1 illustrates, ignoring patent applications can create bias by falsely declaring a patent as completely disruptive. The channels are the same as when backwards citations are truncated in time. Formally, it affects the CD index via $b_{it}$ and $N$ in equation (1), but the driving force is now missing citations to patent applications. As shown by equations (2)-(6), this can create a bias in the CD index similar to left-truncated backward citations.
Figure B.2: Backward Citations to Patent Application and Corresponding Patent Grant Affects the CD Index
@@
Second, ignoring patent applications can create bias if a patent application is cited in one instance and the subsequent patent grant is cited in another instance. This can occur from citations to predecessor patents or to the focal patent—i.e., applications and grants. Fig. B.2 illustrates this possibility in more detail. Formally, it affects the CD index in equation (1) as before in $b_{it}$ or $N$ but now also in $f_{it}$. This may result in an upward or downward bias of the CD index. An upward bias occurs, for example, when the focal patent cites a patent application and a successor patent cites the focal patent and the patent grant (instead of the patent application). In such a case, the citation to the patent application is missed. A downward bias occurs, for example, if a successor patent only cites the patent application of the focal patent and no predecessor patents (applications or grants) of the focal patent. As this would not be considered a forward citation to the focal patent, it thereby loses some of the disruptive value of the focal patent. The Inventor Protection Act of 1999 may therefore bias the CD index if citations to patent applications for patents published after November 29, 2000, are not properly taken into account.
@@ B.1 Figures
Figure B.3: Examples of Backward Citations in Patent Documents
@@
Figure B.4: Average Backward Citation Age
@@
The figure shows the average age of backward citations for various aggregate WIPO technologies, as determined in Table B.3. The left panel (Truncation Methodology) shows that the average backward citation ages increase in each technology because citations before 1976 are not considered. The middle panel (No Truncation Methodology) shows average backward citation ages are relatively stable in each technology up to 2005 and then increase markedly post-2005 because citations to patent applications are not considered. The right panel (No Truncation + Patent Application Citation Methodology) shows average backward citation ages are relatively stable in each technology and over time.
Figure B.5: Average CD Index By Various Methods
@@
The figure shows the average CD index calculated using four different methodologies. Methodology (I) via Park et al. (2023) excludes backward citations to patent grants before 1976 and backward citations to patent applications. Methodology (II) includes backward citations to patent grants before 1976, but excludes backward citations to patent applications. Methodology (III) includes backward citations to patent grants before 1976 and to patent applications that have received patent grants by the end of 2021. This methodology replaces the application number with the corresponding patent grant number and publication date of the application with the publication date of the patent grant. Finally, methodology (IV) includes backward citations to patent grants before 1976, to patent applications that have received patent grants by the end of 2021, and to patent applications that have not received patent grants by the end of 2021. This methodology replaces the application numbers with their corresponding granted patent numbers while keeping the original application dates, which is necessary to ensure that all citations are taken into account as illustrated by Fig. B.2. Note methodology (IV) lies slightly below methodology (III) after 2005 due to the inclusion of backward citations to patent applications that did not receive patent grants and thus would not appear prior to the Act’s passage. Citations to patent applications can increase the number of citations to predecessor patents, which in turn may lead to a decrease in the CD index as the value of $N$ in equation (1) is larger. In this case, however, the decline in the CD index does not necessarily suggest a decline in the number of disruptive innovations. Instead, it could be due to what Petersen et al. (2019) terms "citation inflation".
@@ B.2 Tables
Table B.1: Truncated vs. Non-Truncated Distribution of Patents’ $CD_{5}$ Index for 1980
@@
    The table shows how the $CD_{5}$ values of patents of a particular group calculated using the Park et al. (2023) methodology are distributed across the $CD_{5}$ values using the correct methodology without truncation. Reading example: The last row of the last column shows that only 15 percent of all patents with a $CD_{5}$ index between 0.75 and 1 and calculated by the Park et al. (2023) methodology actually belong to the same group when all backward citations are taken into account.
Table B.2: Truncated vs. Non-Truncated Distribution of Patents’ $CD_{5}$ Index for 2016
@@
    The table shows how patents belonging to a particular group of $CD_{5}$ values calculated with the Park et al. (2023) methodology are distributed across the $CD_{5}$ values using the correct methodology without truncation and considering citations to patent applications. Reading example: The last row of the last column shows that only 12 percent of all patents with a $CD_{5}$ index between 0.75 and 1 calculated by the Park et al. (2023) methodology belong to the same group when all backward citations and citations to patent applications are taken into account.
Table B.3: Aggregation of Technologies

### v2

@@ B CD Index and Patent Law Change
Abbildung B.1: Excluding Backward Citations to Patent Applications Affects the CD Index
@@
First, as Fig. B.1 illustrates, ignoring citations to patent applications can create a bias by falsely declaring a patent as completely disruptive. The channels are the same as when backwards citations are truncated in time. Formally, it affects the CD index via $b_{it}$ and $N$ in equation (1), but the driving force is now missing citations to patent applications. As shown by equations (2)-(6), this can create a bias in the CD index similar to left-truncated backward citations.
Abbildung B.2: Backward Citations to Patent Application and Corresponding Patent Grant Affects the CD Index
@@
Second, ignoring citations to patent applications can create a bias if a patent application is cited in one instance and the subsequent patent grant is cited in another instance. This can occur from citations to predecessor patents or to the focal patent—i.e., applications and grants. Fig. B.2 illustrates this possibility in more detail. Formally, it affects the CD index in equation (1) as before in $b_{it}$ or $N$ but now also in $f_{it}$. This may result in an upward or downward bias of the CD index. An upward bias occurs, for example, when the focal patent cites a patent application and a successor patent cites the focal patent and the patent grant (instead of the patent application). In such a case, the citation to the patent application is missed. A downward bias occurs, for example, if a successor patent only cites the patent application of the focal patent and no predecessor patents (applications or grants) of the focal patent. As this would not be considered a forward citation to the focal patent, it thereby loses some of the disruptive value of the focal patent. The Inventor Protection Act of 1999 may therefore bias the CD index if citations to patent applications for patents published after November 29, 2000, are not properly taken into account.
@@ B.1 Figures
Abbildung B.3: Examples of Backward Citations in Patent Documents
@@
Abbildung B.4: Average Backward Citation Age
@@
The figure shows the average age of backward citations for various aggregate WIPO technologies, as determined in Table B.3. The vertical line at 2010 marks the end of the Park et al. (2023) analysis. The left panel (Truncation Methodology) shows that the average backward citation ages increase in each technology because citations before 1976 are not considered. The middle panel (No Truncation Methodology) shows average backward citation ages are relatively stable in each technology up to 2005 and then increase markedly post-2005 because citations to patent applications are not considered. The right panel (No Truncation + Patent Application Citation Methodology) shows average backward citation ages are relatively stable in each technology and over time.
Abbildung B.5: Average CD Index By Various Methods
@@
The figure shows the average CD index calculated using four different methodologies. The vertical line at 2010 marks the end of the Park et al. (2023) analysis. Methodology (I) via Park et al. (2023) excludes backward citations to patent grants before 1976 and backward citations to patent applications. Methodology (II) includes backward citations to patent grants before 1976, but excludes backward citations to patent applications. Methodology (III) includes backward citations to patent grants before 1976 and to patent applications that have received patent grants by the end of 2021. This methodology replaces the application number with the corresponding patent grant number and publication date of the application with the publication date of the patent grant. Finally, methodology (IV) includes backward citations to patent grants before 1976, to patent applications that have received patent grants by the end of 2021, and to patent applications that have not received patent grants by the end of 2021. This methodology replaces the application numbers with their corresponding granted patent numbers while keeping the original application dates, which is necessary to ensure that all citations are taken into account as illustrated by Fig. B.2. Note methodology (IV) lies slightly below methodology (III) after 2005 due to the inclusion of backward citations to patent applications that did not receive patent grants and thus would not appear prior to the Act’s passage. Citations to patent applications can increase the number of citations to predecessor patents, which in turn may lead to a decrease in the CD index as the value of $N$ in equation (1) is larger. In this case, however, the decline in the CD index does not necessarily suggest a decline in the number of disruptive innovations. Instead, it could be due to what Petersen et al. (2019) terms "citation inflation".
Abbildung B.6: Correlation Over Time Between Different Variables
The figure shows correlations between different methods over time for (1) backward citations of focal patents, (2) citations to focal patents (5 years after publication of the focal patent), (3) citations to focal patents and/or their predecessor patents (5 years after publication of the focal patent), (4) and the $CD_{5}$ index. The blue lines show the correlations between Methodology I and II (Park et al., 2023, Methodology and No Truncation Methodology), the orange line between I and III (Park et al., 2023, Methodology and No Truncation + Patent Application Citation Methodology), and the green line between II and III (No Truncation Methodology and No Truncation + Patent Application Citation Methodology). The vertical line at 2010 marks the end of the Park et al. (2023) analysis.
Abbildung B.7: Number of Very Highly Disruptive Patents Across Technologies
The figure illustrates the number of very highly disruptive patents for various aggregate WIPO technologies, as determined in Appendix Table B.3. The vertical line at 2010 marks the end of the Park et al. (2023) analysis. The respective horizontal lines show the values for 1980 for comparison purposes. The blue line shows the number of patents when backward citations to granted U.S. patents published before 1976 are excluded, per the Park et al. (2023) methodology. The green line shows the number of patents when all backward citations to patent grants and applications are included.
Abbildung B.8: CD Index With Varying Observation Periods For Forward Citations
The figure shows the average $CD$ Index for different observation periods of forward citations calculated using the Park et al. (2023) methodology and our methodology that includes all backward citations to patent grants and patent applications. The vertical line at 2010 marks the end of the Park et al. (2023) analysis. The left-hand panel considers forward citations until 10 years after the publication year of a focal patent. Since data are used until the end of 2021, this value can only be calculated until 2011. The right-hand panel considers all citations until the end of 2021.
Abbildung B.9: Different Variants of the CD Index
The figure shows different variants of the average $CD$ index as used by Park et al. (2023), calculated using the Park et al. (2023) methodology and our methodology that includes all backward citations to patent grants and patent applications. Each vertical line at 2010 marks the end of the Park et al. (2023) analysis. The first variant, shown in the upper left panel, normalizes the $CD_{5}$ index by the number of patents that cite only the predecessors of the focal patents. In particular, the denominator of the CD index (equation 1) is adjusted by subtracting the number of backward citations of the focal patent: $N_{norm}=N_{F}+N_{B}+N_{R}-N_{b}$, where $N_{F}$ is the sum of all citations made only to the focal patent and not to any of its predecessors patents, $N_{B}$ all citations to the focal patent and parts of its predecessors, and $N_{R}$ all citations to parts of its predecessors only. In cases where $N_{R}-N_{b}<0$, we set $N_{R}-N_{b}=0$ as in Park et al. (2023). The second variant, shown in the upper right panel, normalizes the $CD_{5}$ index by the average number of backward citations of all patents of the corresponding two-digit WIPO technology and publication year: $N_{norm}^{mean}=N_{F}+N_{B}+N_{R}-N_{b}^{mean}$. In cases where $N_{R}-N_{b}^{mean}<0$, we set $N_{R}-N_{b}^{mean}=0$ as in Park et al. (2023). The third variant, shown in the lower left panel, is a $CD_{5}$ index proposed by Leydesdorff et al. (2021) and considers in the nominator only forward citations made to the focal patent but not to its predecessors: $DI^{*}=\frac{N_{F}}{N_{F}+N_{B}+N_{R}}$. The fourth variant, shown in the lower right panel, is a $CD_{5}$ index proposed by Bornmann et al. (2020) and includes in the denominator only forward citations made solely to the focal patent and to the focal patent and parts of its predecessors: $DI_{1}^{no\penalty 10000\ k}=\frac{N_{F}-N_{B}}{N_{F}+N_{B}}$.
@@ B.2 Tables
Tabelle B.1: Truncated vs. Non-Truncated Distribution of Patents’ $CD_{5}$ Index for 1980
@@
    The table shows how the $CD_{5}$ values of patents of a particular group calculated using the Park et al. (2023) methodology are distributed across the $CD_{5}$ values using the methodology without truncation. Reading example: The last row of the last column shows that only 15 percent of all patents with a $CD_{5}$ index between 0.75 and 1 and calculated by the Park et al. (2023) methodology actually belong to the same group when all backward citations are taken into account.
Tabelle B.2: Truncated vs. Non-Truncated Distribution of Patents’ $CD_{5}$ Index for 2016
@@
    The table shows how patents belonging to a particular group of $CD_{5}$ values calculated with the Park et al. (2023) methodology are distributed across the $CD_{5}$ values using the methodology without truncation and considering citations to patent applications. Reading example: The last row of the last column shows that only 12 percent of all patents with a $CD_{5}$ index between 0.75 and 1 calculated by the Park et al. (2023) methodology belong to the same group when all backward citations and citations to patent applications are taken into account.
Tabelle B.3: Aggregation of Technologies
@@ 35 Civil engineering other
Literatur
  • Altmetric (2023) Overview of attention for article Papers and patents are becoming less disruptive over time published in nature, january 2023. Note: https://nature.altmetric.com/details/140824206/citationsAccessed: 2023-04-26 Cited by: §1.
  • N. Bloom, C. I. Jones, J. V. Reenen, and M. Webb (2020) Are ideas getting harder to find?. American Economic Review 110 (4), pp. 1104–1144. Cited by: §1, footnote 1.
  • P. Boeing and P. Hünermund (2020) A global decline in research productivity? evidence from china and germany. Economics Letters 197, pp. 109646. Cited by: §1.
  • L. Bornmann, S. Devarakonda, A. Tekles, and G. Chacko (2020) Are disruption index indicators convergently valid? the comparison of several indicator variants with assessments by peers. Quantitative Science Studies 1 (3), pp. 1242–1259. Cited by: §1, Abbildung B.9, footnote 17, footnote 3, footnote 9.
  • S. Braguinsky, J. Choi, Y. Ding, K. Jo, and S. Kim (2023) Mega firms and recent trends in the us innovation: empirical evidence from the us patent data. Note: NBER Working Paper 3146 Cited by: footnote 2.
  • J. S. Chu and J. A. Evans (2021) Slowed canonical progress in large fields of science. Proceedings of the National Academy of Sciences 118 (41), pp. e2021636118. Cited by: §1.
  • Financial Times (2023) Science is losing its ability to disrupt (2023). Vol. 18.1.2023. Note: https://www.ft.com/content/c8bfd3da-bf9d-4f9b-ab98-e9677f109e6d Cited by: §1.
  • R. J. Funk and J. Owen-Smith (2017) A dynamic network measure of technological change. Management science 63 (3), pp. 791–817. Cited by: §1, §A, §2, footnote 3.
  • D. K. Johnson and D. Popp (2003) Forced out of the closet: the impact of the american inventors protection act on the timing of patent disclosure. RAND Journal of Economics 34 (1), pp. 96–112. Cited by: §5.
  • M. Kaltenberg, A. B. Jaffe, and M. E. Lachman (2023) Invention and the life course: age differences in patenting. Research Policy 52 (1), pp. 104629. Cited by: footnote 3.
  • A. Kalyani (2022) The creativity decline: evidence from us patents. Note: Available at SSRN 4318158 Cited by: footnote 2.
  • J. Kuhn, K. Younge, and A. Marco (2020) Patent citations reexamined. The RAND Journal of Economics 51 (1), pp. 109–132. Cited by: §B.
  • C. Leibel and L. Bornmann (2023) What do we know about the disruption index in scientometrics? an overview of the literature. Scientometrics, pp. 1–39. Cited by: §1.
  • L. Leydesdorff, A. Tekles, and L. Bornmann (2021) A proposal to revise the disruption index. Profesional de la información (EPI) 30 (1). Cited by: Abbildung B.9, footnote 17.
  • Y. Lin, C. B. Frey, and L. Wu (2023) Remote collaboration fuses fewer breakthrough ideas. Nature 623 (7989), pp. 987–991. External Links: ISSN 1476-4687 Cited by: footnote 3.
  • M. Park, E. Leahey, and R. J. Funk (2023) Papers and patents are becoming less disruptive over time. Nature 613 (7942), pp. 138–144. Cited by: §1, §1, §1, §1, §1, §1, §1, §1, §A, §A, Abbildung 1, Abbildung 1, Abbildung 1, Abbildung B.4, Abbildung B.5, Abbildung B.6, Abbildung B.7, Abbildung B.8, Abbildung B.9, item 1, item 1, §2, §2, §2, §2, §2, §3, §3, Abbildung 3, Abbildung 4, §4, §4, Abbildung 5, Abbildung 5, Abbildung 5, §5, §5, Abbildung 6, §6, §6, §6, §6, §7, §7, §7, footnote 13, footnote 15, footnote 16, footnote 17, footnote 4, footnote 7, Is There a Secular Decline in Disruptive Patents? Correcting for Measurement Bias.
  • A. M. Petersen, F. Arroyave, and F. Pammolli (2023) The disruption index is biased by citation inflation. Note: arXiv preprint arXiv:2306.01949 Cited by: §1.
  • A. M. Petersen, R. K. Pan, F. Pammolli, and S. Fortunato (2019) Methods to account for citation inflation in research evaluation. Research Policy 48 (7), pp. 1855–1865. Cited by: Abbildung B.5.
  • X. Ruan, D. Lyu, K. Gong, Y. Cheng, and J. Li (2021) Rethinking the disruption index as a measure of scientific and technological advances. Technological Forecasting and Social Change 172, pp. 121071. External Links: ISSN 0040-1625 Cited by: §1.
  • R. Rust (2023) Rebooting disruptive science: exploring the challenges and potential solutions. Wiley Online Library. Note: European Journal of Clinical Investigation Cited by: §1.
  • U. Schmoch (2008) Concept of a technology classification for country comparisons. Technical report WIPO. Cited by: §A, item 1.
  • The Economist (2023) Papers and patents are becoming less disruptive (2023). Vol. 4.1.2023. Note: https://www.economist.com/science-and-technology/2023/01/04/papers-and-patents-are-becoming-less-disruptive Cited by: §1.
  • The New York Times (2023) What happened to all of science’s big breakthroughs? (2023). Vol. 17.1.2023. Note: https://www.nytimes.com/2023/01/17/science/science-breakthroughs-disruption.html Cited by: §1.
  • L. Wu, D. Wang, and J. A. Evans (2019) Large teams develop and small teams disrupt science and technology. Nature 566 (7744), pp. 378–382. Cited by: footnote 3, footnote 9.
  • I. Yanai and M. J. Lercher (2023) Make science disruptive again. Nature Biotechnology 41, pp. 1–2. Cited by: §1.